Mutiply and simplify:
(4-3 square root 2) (5+2 square root 2)

Answer by stanbon(75887)

(4-3 square root 2) (5+2 square root 2)
----------------------------
Use FOIL:
F: 4*5=20
O: 4*2sqrt2 = 8sqrt(2)
I: 5*-3sqrt(2) = -15sqrt(2)
L: -3sqrt2 *2sqrt2 = -6*2 = -12
---------------------------------
Add them up to get:
8-7sqrt2
==========
Cheers,
Stan H.